hey, help me in my task please, thank you?

A construction crew wants to hoist a heavy beam so that it is standing up straight. They
tie a rope to the beam, secure the base, and pull the rope through a pulley to raise one end
of the beam from the ground. When the beam makes an angle of 40o with the ground, the
top of the beam is 8 ft above the ground.
The construction site has some telephone wires crossing it.. The workers are concerned
that the beam may hit the wires. When the beam makes an angle of 60o with the ground,
the wires are 2 ft above the top of the beam. Will the beam clear the wires on its way to
standing up straight? Explain your answer.

Will the beam clear the wires on its way to standing up straights?

Answer by mananth(16946)

The beam take a circular path and length of beam as the radius.
In triangle ADG
AD = 8/sin 40
= 12.45 ft.
In triangle AMB
sin 60 = MB / AM
MB = 12.45 * sin 60
MB = 10.78
The cables are two ft above
10.78+2 = 12.78 ft
The length of ladder is 12.45
the beam clear the wires on its way to
standing up straight.
.